A key falls from a bridge that is 38 m above the water. It falls directly into a model boat, moving with constant velocity, that was 11 m from the point of impact when the key was released. What is the speed of the boat?
___ m/s
Answer #1

using the follmula y - y= 0 +Lat? __19+² (U=o) (a=-g) -049 t we have AY - 38 = 9.8 Substitute this value in equation 80 t =